### v1

# CustomVideo: Customizing Text-to-Video Generation with Multiple Subjects

Abstract
Customized text-to-video generation aims to generate high-quality videos guided by text prompts and subject references. Current approaches designed for single subjects suffer from tackling multiple subjects, which is a more challenging and practical scenario. In this work, we aim to promote multi-subject guided text-to-video customization. We propose CustomVideo, a novel framework that can generate identity-preserving videos with the guidance of multiple subjects. To be specific, firstly, we encourage the co-occurrence of multiple subjects via composing them in a single image. Further, upon a basic text-to-video diffusion model, we design a simple yet effective attention control strategy to disentangle different subjects in the latent space of diffusion model. Moreover, to help the model focus on the specific object area, we segment the object from given reference images and provide a corresponding object mask for attention learning. Also, we collect a multi-subject text-to-video generation dataset as a comprehensive benchmark, with 69 individual subjects and 57 meaningful pairs. Extensive qualitative, quantitative, and user study results demonstrate the superiority of our method, compared with the previous state-of-the-art approaches.
Figure 1: Customized text-to-video generation results of our proposed CustomVideo with given multiple subjects (left) and text prompts (below). Our approach can disentangle highly similar subjects, e.g., cat v.s. dog, preserving the fidelity of subjects and smooth motions.
1 Introduction
Text-to-video (T2V) generation Wang et al. (2023a); Zeng et al. (2023); Blattmann et al. (2023a); Kondratyuk et al. (2023) has achieved fantastic progress taking advantages of diffusion models Ho et al. (2020); Song et al. (2020); Lu et al. (2022). Recently, artists have dreamed of generating videos with their own belongings, e.g., pets, which directs a new research direction named customized T2V generation. Although existing methods Zhao et al. (2023); Wei et al. (2023); Jiang et al. (2023) have been proposed to generate videos from a single object, tackling multiple objects still remains a difficult scenario. The key challenge is to ensure the co-occurrence of multiple objects in the generated video and retain their corresponding identities.
A recent work, VideoDreamer Chen et al. (2023b), proposes disen-mix finetuning and human-in-the-loop re-finetuning strategies based on Stable Diffusion Rombach et al. (2022), aiming to generate videos from multiple subjects. However, VideoDreamer falls short in guaranteeing the co-occurrence of multiple subjects and disentangling similar subjects due to its inconsistent object mixing strategy. As shown in Figure 3, the cat can not be consistently generated across all frames during inference. Moreover, VideoDreamer can not tackle a scenario with a foreground object with a background scene, such as car and barn (see Figure 3).
In contrast, in our approach, we ensure the co-occurrence of multiple objects during model training, which encourages the model to capture the presence of different subjects simultaneously, thereby facilitating co-occurrence during inference. Additionally, we propose an attention control mechanism to disentangle multiple subjects during training, effectively guiding the model to focus on the corresponding subject area while disregarding irrelevant parts of the image. To facilitate this process, we incorporate a ground truth object mask, obtained through segmentation either from a model like SAM Kirillov et al. (2023) or provided by human annotators, as supervision during optimization. Our attention mechanism consists of two key designs that contribute to the disentanglement of subjects. Firstly, we highlight the corresponding subject area using the ground truth object mask on the cross-attention map, aligning learnable text tokens with subject identity. Secondly, we optimize the cross-attention map towards a slight negative value, excluding the desired subject and mitigating the influence of irrelevant areas in the input image. To comprehensively evaluate our proposed approach, we have curated a diverse dataset covering a wide range of categories including pet, person, toy, transportation, scene, plant, wearable item, and so on. Beyond the MultiStudioBench Chen et al. (2023b) which includes 12 subject pairs, our dataset CustomStudio comprises 69 individual subjects and 57 subject pairs, notably featuring some challenging scenarios involving visually similar objects. Through extensive experiments on this benchmark dataset, we provide qualitative, quantitative, and user study results that demonstrate the superiority of our method in generating high-quality videos with customized subjects. In summary, our contributions are as follows:
  • We propose CustomVideo, a novel multi-subject driven T2V generation framework, powered by a simple yet effective co-occurrence and attention control mechanism.
  • We collect a multi-subject T2V dataset and build it as a comprehensive benchmark. We highlight that our benchmark covers a wide range of subject categories and diverse subject pairs over them.
  • Our method consistently outperforms previous state-of-the-art approaches in terms of diverse metrics, including Textual Alignment, Image Aligenment, and Temporal Consistency. More critically, in the user study, we obtain significant improvement with the scores 3×3$\times$3 × higher than these methods.
Figure 2: The overview of our proposed CustomVideo. We propose a simple yet effective co-occurrence and attention control mechanism with mask guidance to preserve the the fidelity of subjects for multi-subject driven text-to-video generation. During the training stage, only the key and value weights in the cross attention layers are fine-tuned. In the inference stage, given a text prompt integrated with learned text token, we can easily obtain high-quality videos with specific subjects.
2 Related Work
2.1 Text-to-Video Generation
Text-to-video generation has made significant advancements in recent years Chen et al. (2023a); Duan et al. (2023); Wang et al. (2023a). Early approaches employed GANs Balaji et al. (2019); Skorokhodov et al. (2022) and VQVAE Hong et al. (2022); Villegas et al. (2023), while more recent works have explored diffusion models to generate high-quality videos He et al. (2022); Wang et al. (2023c); Zhang et al. (2023a, b). Make-A-Video Singer et al. (2023) utilizes a pre-trained image diffusion model with additional temporal attention finetuning. VideoLDM Blattmann et al. (2023b) introduces a multi-stage alignment approach in the latent space to generate high-resolution and temporally consistent videos. Other methods Khachatryan et al. (2023); Guo et al. (2023) generate videos with an image as the first frame and randomly initialized subsequent frames. To enhance controllability, VideoComposer Wang et al. (2023b) incorporates additional guidance signals, such as depth maps, to produce desired videos alongside text inputs. Tune-A-Video Wu et al. (2023) proposes a temporal self-attention module that fine-tunes a pre-trained image diffusion model, achieving successful generation of videos with specific text guidance. Furthermore, diffusion-based video-to-video editing approaches Qi et al. (2023); Geyer et al. (2023) have also been proposed for practical usages.
2.2 Subject-driven Customization
There has been a growing interest in customizing pre-trained image and video diffusion models for personalized generation. Customization involves generating images and videos with specific subjects, typically based on a few reference images. For image diffusion customization, Textual Inversion Gal et al. (2023) represents a specific object as a learnable text token using only a few reference images. This learned text token can then be integrated into a sentence to generate personalized images during the inference stage. Additionally, DreamBooth Ruiz et al. (2023) fine-tunes the weights of the diffusion model to improve fidelity in image generation. Several works Kumari et al. (2023); Liu et al. (2023a, b) have explored personalized image diffusion with multiple subjects, focusing on parameter-efficient finetuning and text embedding learning. While there have been initial attempts to customize video diffusion, such as VideoAssembler Zhao et al. (2023) and VideoBooth Jiang et al. (2023), which use reference images to personalize the video diffusion model while preserving subject identity, and DreamVideo Wei et al. (2023), which decouples the learning process for subject and motion customization, these methods are designed for single objects and cannot handle multiple subjects when given. A recent work, VideoDreamer Chen et al. (2023b), proposes multi-subject driven video customization through disen-mix finetuning strategy with LoRA Hu et al. (2022). However, the generated videos do not guarantee the co-occurrence of multiple subjects, or disentanglement of different subjects. In this work, we propose a simple yet effective co-occurrence and attention mechanism that disentangles multiple subjects using masks as guidance while preserving the co-occurrence of subjects in the generated videos.
3 Method
3.1 Preliminary: Text-to-Video
Video diffusion models (VDMs) Wang et al. (2023a); Duan et al. (2023); Chen et al. (2023a) generate videos by gradually denoising a randomly sampled Gaussian noise $\epsilon$, following an iterative denoising process that resembles a reverse procedure of a fixed-length Markov Chain. This iterative denoising allows VDMs to capture the temporal dependencies presented in video data. To be specific, a video diffusion model $\theta$ predicts the added noise at each timestep $t$ given a text condition $c$, where $t\in\{1,2,\cdots,T\}$. Thus, the training objective for this process can be expressed as a reconstruction loss:
$\mathcal{L}_{recon}=\mathbb{E}_{\epsilon,\boldsymbol{z},c,t}\left[\left\|\epsilon-\epsilon_{\theta}\left(\boldsymbol{z}_{t},\mathcal{T}(c),t\right)\right\|_{2}^{2}\right],$ (1)
where $\boldsymbol{z}\in\mathbb{R}^{B\times L\times H\times W\times D}$ is the latent code of the input video with batch size $B$, video length $L$, height $H$, width $W$, and latent dimension $D$. $\epsilon_{\theta}$ is the noise prediction from the model. $\mathcal{T}$ is a pre-trained text encoder. $\boldsymbol{z}_{t}$ is obtained by adding noise to the ground truth $\boldsymbol{z}_{0}$ with $\boldsymbol{z}_{t}=\alpha_{t}\boldsymbol{z}_{0}+\sqrt{1-\alpha_{t}^{2}}\epsilon$, where $\alpha_{t}$ is a diffusion hyperparameter. In this work, we utilize zeroscope [^1] https://huggingface.co/cerspense/zeroscope_v2_576w T2V model as our base model, which is built upon a 3D UNet, with spatial and temporal modeling for generating high-quality videos.
3.2 CustomVideo with Multiple Subjects
Figure 2 provides an overview of our proposed CustomVideo framework, which enables multi-subject driven T2V customization. During the training stage, we employ a concatenation technique to ensure the co-occurrence of multiple subjects. Specifically, we combine these subjects into a single image, facilitating the learning of multi-subject patterns by the model. To address the challenge of entanglement among highly similar subjects, we propose an attention control mechanism. This mechanism ensures that the learnable word tokens align with their corresponding regions on the cross-attention map. By achieving this alignment, we enable the model to disentangle different subjects and enhance the quality of the generated videos. The training process focuses on training the subject-related learnable word tokens, as well as the key and value weights within the cross-attention layers of the UNet architecture, adopting a parameter-efficient finetuning approach. During the inference stage, users only need to provide a text prompt description integrated with the corresponding learned word tokens to generate high-quality videos aligned with their preferences. In the following, we will delve into the detailed workings of CustomVideo.
Co-occurrence Control.
In the context of multi-subject driven T2V generation, ensuring that the model consistently learns to generate videos with multiple subjects is crucial. In a previous work, VideoDreamer Chen et al. (2023b), the authors proposed a disen-mix strategy for multi-subject driven generation by fine-tuning the model using both single subject images and concatenated images of multiple subjects. However, we have identified that this mixing strategy can confuse the model due to the inconsistent number of subjects present in a single image. Consequently, it leads to unstable generation of videos with multiple subjects (see line ‘both single and concat’ in Figure 5). In our approach, we have found that fine-tuning the model using only concatenated images of multiple subjects is sufficient to ensure the co-occurrence of multiple subjects in the generated videos. Additionally, we have observed that providing clear subjects without background aids the model in learning the specific characteristics of the subjects more effectively. To achieve this, we perform background removal on the subject images. This can be accomplished manually or by employing an automatic tool such as the SAM model Kirillov et al. (2023).
Attention Control.
Ensuring the co-occurrence of multiple subjects in generated videos is achieved through image concatenation. However, a more challenging task is preserving the distinct characteristics of each subject without interference. Simply fine-tuning the model with concatenated images can lead to confusion among subject characteristics. As shown in Figure 5 (line ‘w/o pos. attn.’), the generated cat predominantly resembles the shape of the provided dog sample, despite having a similar color texture to the provided cat. Thus, disentangling the multiple subjects becomes crucial to generate high-quality videos that faithfully represent each subject’s characteristics.
[…]
3.3 Model Training and Inference
Training Strategy.
During training, we employ a parameter efficient fine-tuning strategy by only making the weights of key and value in all of the cross attention layers trainable. Inspired by previous text-to-image personalization Ruiz et al. (2023); Kumari et al. (2023), we conduct class-specific prior preservation to improve the diversity of generated videos and alleviate the issue of language drift. The loss for prior preservation is formulated as the following:
$\mathcal{L}_{recon}^{pr}=\mathbb{E}_{\epsilon^{\prime},\boldsymbol{z}^{\prime},c^{\prime},t^{\prime}}\left[\left\|\epsilon^{\prime}-\epsilon_{\theta}^{\prime}\left(\boldsymbol{z}_{t^{\prime}}^{\prime},\mathcal{T}(c^{\prime}),t^{\prime}\right)\right\|_{2}^{2}\right],$ (3)
where $\boldsymbol{z}^{\prime}$ is the latent code of the input class image. $\epsilon_{\theta}^{\prime}$ is the noise prediction from the model $\theta$ and $\epsilon^{\prime}$ is the randomly sampled Gaussian noise. $c^{\prime}$ is the text condition for class image and $t^{\prime}$ is the sampling timestep. To this end, we train our CustomVideo via an end-to-end manner, with the following overall training objective:
$\mathcal{L}=\mathcal{L}_{recon}+\alpha\cdot\mathcal{L}_{attn}+\beta\cdot\mathcal{L}_{recon}^{pr},$ (4)
where $\alpha$ and $\beta$ are two hyper-parameters to control the weight of attention control and prior preservation, respectively.
Inference.
During inference, CustomVideo only requires a specific text prompt with corresponding learned word token integrated to generate a required video. To be note that ground truth masks of subjects are not required in inference. With only text prompt as input, our proposed CustomVideo can generate videos with high diversity while not losing the the fidelity of subjects and motion smoothness.
4 Experiment
Figure 3: Qualitative results of our CustomVideo with comparison to SOTA methods, including DreamBooth, CustomDiffusion, and VideoDreamer. The first line indicates the given subjects, while each line indicates the frames generated by the corresponding method. The bottom line is the text prompt used during inference. We can observe that our CustomVideo can generate videos with much better fidelity of subjects compared with previous SOTA methods.
4.1 Experimental Setup
Dataset.
We collect a dataset CustomStudio with 69 individual subjects and construct 57 meaningful pairs for multi-subject driven T2V generation. These subjects are adapted from DreamBooth Ruiz et al. (2023), CustomDiffusion Kumari et al. (2023) and internet. The objects cover a wide range of categories, including pet, person, toy, transportation, scene, plant, wearable item, and so on. Each pair of subjects has 10 different text prompts, which are designed with different contexts, actions, and so on.
Implementation Details.
We train CustomVideo for 500 steps with AdamW Loshchilov and Hutter (2017) optimizer (batch size 2, learning rate 4e-5, and weight decay 1e-2). For class-specific prior preservation, we collect 200 class images from LAION-400M Schuhmann et al. (2021) for each subject. Note that the corresponding class images are also concatenated during training phase. The negative value $\eta$ in the mask $\boldsymbol{\mathcal{M}}$ is set as -1e-8. The weight parameters $\alpha$ and $\beta$ in Eq. (4) are set as 0.1 and 1.0, respectively. During inference, we perform 50 steps denoising with DDIM Song et al. (2020) sampler and classifier-free guidance Ho and Salimans (2022). The resolution of generated 24-frame videos are $576\times 320$ with 8 fps. Interestingly, we find that the trained weights and word tokens learned from low resolution zeroscope T2V model can be directly loaded for high resolution one [^2] https://huggingface.co/cerspense/zeroscope_v2_XL to generate personalized videos with $1024\times 576$ resolution, in which no additional training computation cost is required. Our CustomVideo is implemented based on Diffusers von Platen et al. (2022). The training phase takes about 8 minutes for a subject pair on 1 RTX 3090 GPU. Meanwhile, it takes about 1 minute and 3 minutes to generate a low and high resolution video on 1 RTX 3090 GPU, respectively.
Comparison Methods.
Except for VideoDreamer Chen et al. (2023b), we also consider adapting previous SOTA image-based multi-subject driven methods to a video scenario for comparison, including DreamBooth Ruiz et al. (2023) and CustomDiffusion Kumari et al. (2023). DreamBooth trains all of the weights in a diffusion model to personalize the generation from multiple subjects. CustomDiffusion fine-tune the diffusion model with joint training on multiple subjects. VideoDreamer proposes a disen-mix and human-in-the-loop fine-tuning strategy for multi-subject customization of T2V model. For fair comparison, we implement all the comparison method and our CustomVideo based on zeroscope model.
Table 1: Quantitative results of our CustomVideo with comparison to SOTA methods, including DreamBooth, CustomDiffusion, and VideoDreamer. Our proposed CustomVideo consistently outperforms previous SOTA methods for all 4 evaluation metrics.
Method CLIP-T $\uparrow$ CLIP-I $\uparrow$ DINO-I $\uparrow$ T. Cons. $\uparrow$
DreamBooth 0.6476 0.6015 0.3098 0.7209
CustomDiffusion 0.6581 0.6218 0.3186 0.7548
VideoDreamer 0.6649 0.6314 0.3471 0.7362
CustomVideo (ours) 0.7051 0.6749 0.3955 0.8142
Evaluation Metrics.
Following previous works Wei et al. (2023); Chen et al. (2023b), we quantitatively evaluate our CustomVideo with the following 4 metrics: 1) CLIP Textual Alignment computes the average cosine similarity between the generated frames and text prompt with CLIP Radford et al. (2021) ViT-B/32 Dosovitskiy et al. (2021) image and text models; 2) CLIP Image Alignment calculates the average cosine similarity between the generated frames and subject images with CLIP ViT-B/32 image model; 3) DINO Image Alignment measures the average visual similarity between generated frames and reference images with DINO Caron et al. (2021) ViT-S/16 model; 4) Temporal Consistency Esser et al. (2023) evaluates the average cosine similarity of between all consecutive frame pairs in the generated videos.
4.2 Main Results
Quantitative Results.
We conduct quantitative experiments on our collected dataset. To ensure a thorough analysis, we generate videos using 10 individual prompts and 4 random seeds for each pair of subjects, resulting in a total of 2,280 generated videos for each method. We then evaluate the quality of the generated videos using four metrics, and the results are presented in Table 1. The table clearly demonstrates that our proposed method is capable of generating videos that are better aligned with the given subjects. This improvement can be attributed to our specially designed co-occurrence and attention control mechanisms, which effectively disentangle and preserve the fidelity of the subjects. Notably, our CustomVideo outperforms the most recent method, VideoDreamer, by 6.05%, 6.89%, and 13.94% in terms of CLIP Textual Alignment, CLIP Image Alignment, and DINO Image Alignment, respectively. Furthermore, our CustomVideo generates videos with significantly higher temporal consistency compared to state-of-the-art methods, as indicated in Table 1. For instance, CustomVideo surpasses VideoDreamer by 10.59% in terms of Temporal Consistency.
Figure 4: User study. Our CustomVideo achieves the best human preference compared with 3 SOTA comparison methods in terms of Textual Alignment, Image Alignment, and Overall Quality.
Figure 5: Qualitative results for component analysis of our proposed CustomVideo. We find that ensuring concatenating subjects during training is effective for guaranteeing the co-occurrence in the generated video. Moreover, our attention mechanism could disentangle different subjects.
Table 2: Quantitative results for component analysis of our proposed CustomVideo. We observe a significant performance drop when removing subjects concatenation or background removal. In practice, we suggest simultaneously using positive and negative attention mechanisms to obtain the best results.
Method CLIP-T $\uparrow$ CLIP-I $\uparrow$ DINO-I $\uparrow$ T. Cons. $\uparrow$
w/o remove bg 0.6676 0.6248 0.3498 0.7891
w/o concat 0.6189 0.6371 0.3572 0.7936
both single and concat 0.6148 0.6319 0.3208 0.7983
w/o pos. attn. 0.6858 0.6402 0.3569 0.8015
w/o neg. attn. 0.6926 0.6392 0.3694 0.8075
CustomVideo (ours) 0.7051 0.6749 0.3955 0.8142
Qualitative Results.
We present the qualitative comparison results in Figure 3. From these results, we observe that our CustomVideo approach effectively ensures the co-occurrence of multiple subjects and successfully disentangles different subjects. However, both DreamBooth and CustomDiffusion fail to capture the structural color information of the car provided in our experiments. The dominant black color from the car window obscures the entire car, resulting in low fidelity. Moreover, the generated frames from VideoDreamer lack consistency, as some frames depict two cars while others only show one car. Additionally, VideoDreamer fails to capture accurate color information, such as the color of the car door. In contrast, our CustomVideo method excels in handling such challenging scenes and foreground subject scenarios, effectively capturing the intricate structural details of the provided car for video generation. Similarly, when considering the case of ‘cat’ and ‘dog’, our approach also demonstrates superior capability in generating high-quality videos.
Human Preference Study.
To further validate our method, we conduct human evaluations on our CustomVideo with comparison to 3 SOTA methods. In this study, we collect 900 answers from 25 independent human raters with the following questions: 1) which one is aligned to the text prompt best? 2) which one is aligned to the subject images best? 3) which one has the best overall quality? The results are shown in Figure 4. It can be found that our proposed CustomVideo is the most preferred one from the human side from all these three dimensional evaluation.
4.3 Ablation Studies
We conduct ablation analysis of our method from 2 aspects: 1) the effect of each component in our method; 2) the effect of hyper-parameters in attention control mechanism.
Table 3: Quantitative results of the weight of attention loss. In practice, $\alpha$=0.1 yields the best result and is adopted in all experiments.
$\alpha$ CLIP-T $\uparrow$ CLIP-I $\uparrow$ DINO-I $\uparrow$ T. Cons. $\uparrow$
1.0 0.6784 0.6651 0.3792 0.8015
0.1 0.7051 0.6749 0.3955 0.8142
0.01 0.6892 0.6584 0.3817 0.8092
Table 4: Quantitative results of the negative value. Interestingly, $\eta$=-1e-8 achieves the best result in terms of all the considered metrics. Practically, we adopt this setting in all the experiments.
$\eta$ CLIP-T $\uparrow$ CLIP-I $\uparrow$ DINO-I $\uparrow$ T. Cons. $\uparrow$
-1e-5 0.6891 0.6582 0.3917 0.7996
-1e-8 0.7051 0.6749 0.3955 0.8142
-1e-11 0.6918 0.6627 0.3895 0.8079
Component Analysis.
We conduct a thorough analysis of each component in our CustomVideo, presenting both qualitative and quantitative results in Figure 5 and Table 2. These findings provide valuable insights into the importance and effectiveness of each component. One crucial observation is the significance of ensuring co-occurrence during the fine-tuning process. We notice a significant drop in performance when multiple subjects are not concatenated into a single image (lines ‘w/o concat’ and ‘CustomVideo’ in Table 2). The absence of subject concatenation leads to the domination of one single subject during fine-tuning, which is deemed unacceptable (line ‘w/o concat’ in Figure 5). Moreover, we conduct an ablation study on fine-tuning the diffusion model with both single subjects and concatenated subjects to examine if the single subject could aid in learning corresponding characteristics. Surprisingly, the results reveal that adding a single subject to the training process proved detrimental, resulting in inconsistent generation during inference (line ‘both single and concat’ in Figure 5). Furthermore, we observe that removing the background from the given subject images significantly improves the generation of high-quality videos. By eliminating the background, the diffusion model can focus solely on learning the characteristics of the given subjects. The comparison between videos generated with and without background removal clearly illustrates the monotonous nature of videos lacking this component (line ‘w/o remove bg’ in Figure 5). Our attention control mechanism plays a crucial role in aligning the generated videos with the given subjects. The positive attention guidance in particular improves the CLIP Image Alignment metric by 5.42%, as demonstrated in Table 2. The negative attention guidance also proved beneficial in promoting better image alignment. Moreover, the attention control mechanism significantly enhances the temporal consistency of the generated videos, a crucial factor in video generation. The comparison between videos generated without positive attention guidance and videos generated without negative attention guidance clearly showcases the positive impact of both components (lines ‘w/o pos. attn.’ and ‘w/o neg. attn.’ in Figure 5). The positive guidance specifically preserves the unique characteristics of a subject, such as distinguishing between a cat and a dog, while the negative guidance weakens the influence of other subjects on a specific subject. By incorporating both positive and negative guidance, our CustomVideo excels in generating videos with high subject fidelity and remarkable temporal consistency.
Figure 6: Failure cases of our CustomVideo. Our CustomVideo may fail to capture some nuanced details of subjects when meeting extremely similar pairs of subjects, such as ‘sloth’ and ‘teddybear’ (first line). Also, our approach fails to generate vivid facial contents with a global view, such as ‘man’ and ‘bike’ (second line)
Effect of Hyper-parameters in Attention Control.
In addition to component analysis, we also conduct investigations into the effects of two important parameters: the weight of attention loss ($\alpha$) and the negative value ($\eta$) used in the guidance mask ($\boldsymbol{\mathcal{M}}$). The quantitative results of these investigations are presented in Table 3 and Table 4. Regarding the weight of attention loss, we observe that the best performance is achieved when $\alpha$ is set to 0.1. This value results in optimal alignment between the generated videos and the given subjects, indicating the importance of appropriately balancing the attention guidance during training. For the negative value in the guidance mask, our findings reveal that even a slight negative value is sufficient to enhance the quality of T2V generation. This implies that incorporating negative attention guidance can effectively suppress the influence of other subjects on a specific subject, improving the generation quality.
4.4 Limitations
In Figure 6, we present some failure cases encountered during our experiments. One challenge arises when generating subjects that are extremely similar, such as a sloth and a teddy bear, where capturing fine-grained details becomes difficult. Additionally, since our method relies on the ability of base model, it would fail if the base model could not generate, such as small faces. To address the failure cases, we can improve CustomVideo by using a stronger T2V base model, like Stable Video Diffusion Blattmann et al. (2023a). Additionally, we can implement mechanisms to prioritize and emphasize nuanced characteristics in the given images during training.
5 Conclusions
This paper provides a novel framework CustomVideo for multi-subject driven T2V generation, powered by a simple yet effective co-occurrence and attention control mechanism. To evaluate the proposed framework, we collect a comprehensive dataset consisting of 69 individual subjects and 57 meaningful pairs. Extensive experiments and human preference study on collected dataset demonstrate the effectiveness of CustomVideo, outperforming existing methods significantly.
Appendix A Additional Ablation Studies
Layers of Cross Attention Maps.
We investigates the effect of how we extract the cross attention maps from the cross attention layers. There are four levels of cross attention layers in the UNet, in which the sizes cross attention maps are different. Taking the resolution $576\times 320$ of our generated low resolution video as an example, the sizes of cross attention maps are $72\times 40$ ($\mathcal{\ell}_{1}$), $36\times 20$ ($\mathcal{\ell}_{2}$), $18\times 10$ ($\mathcal{\ell}_{3}$), and $9\times 5$ ($\mathcal{\ell}_{4}$). Here, we study the effect of different levels of cross attention maps, the results are shown in Table 5. We observe a significant performance drop with too large ($\mathcal{\ell}_{1}$) or small ($\mathcal{\ell}_{4}$) cross attention maps. Regarding the cross attention maps with middle size from $\mathcal{\ell}_{2}$ and $\mathcal{\ell}_{3}$, $\mathcal{\ell}_{3}$ works better. Thus, in our experiments, we extract cross attention maps from $\mathcal{\ell}_{3}$ for attention control.
Appendix B CustomStudio Dataset
We collect a dataset CustomStudio for multi-subject driven video generation. This dataset is composed of 69 individual objects and 57 meaningful pairs. The objects cover a wide range of categories, including decor item, food, furniture, instrument, jewelry, luggage, person, pet, plant, plushie, scene, thing, toy, transportation, and wearable item. The images are adopted from DreamBooth Ruiz et al. (2023), CustomDiffusion Kumari et al. (2023) and internet. An overview of the samples in CustomStudio dataset is shown in Figure 7.
Appendix C More Experimental Details
All of the training process is under fp16 mixed precision with accelerate package Gugger et al. (2022). We use data augmentations during training, including randomly horizontal flip, randomly crop and resize, together with corresponding prompt change (‘very small’ or ‘close up’ appends before the prompt).
DreamBooth.
For efficient fine-tuning, we utilize LoRA Hu et al. (2022) to adapt the UNet under DreamBooth Ruiz et al. (2023). The rank of LoRA module is set as 4. We use AdamW Loshchilov and Hutter (2017) optimizer, with learning rate 4e-4, weight decay 1e-2, batch size 2, and training steps 1000.
CustomDiffusion.
The weights of query and value in all of the cross attention layers of UNet are fine-tuned in CustomDiffusion Kumari et al. (2023). We use AdamW Loshchilov and Hutter (2017) optimizer, with learning rate 4e-5, weight decay 1e-2, batch size 2, and training steps 500.
VideoDreamer.
LoRA Hu et al. (2022) is utilized for fine-tuning the UNet and text encoder in VideoDreamer Chen et al. (2023b). The AdamW Loshchilov and Hutter (2017) optimizer is used with learning rate 5e-5, weight decay 1e-2, batch size 2, and training steps 500.
CustomVideo.
The weights of query and value in all of the cross attention layers of UNet are fine-tuned. The learnable word tokens are initialized with the tokens of the corresponding class names. The sampling step for high resolution video is 30.
Table 5: Quantitative results of cross attention levels. We find that only using cross attention map from $\ell_{3}$ works best.
$\mathcal{\ell}_{1}$ $\mathcal{\ell}_{2}$ $\mathcal{\ell}_{3}$ $\mathcal{\ell}_{4}$ CLIP-T $\uparrow$ CLIP-I $\uparrow$ DINO-I $\uparrow$ T. Cons. $\uparrow$
0.6134 0.5759 0.3023 0.8125
0.6694 0.6239 0.3450 0.7346
0.7051 0.6749 0.3955 0.8142
0.6144 0.6545 0.3426 0.7992
0.6882 0.6177 0.3010 0.8076
0.7047 0.5801 0.3515 0.7930
0.6502 0.6407 0.3651 0.7875
0.6267 0.6734 0.3746 0.7594
0.6869 0.5885 0.3894 0.7605
Figure 7: The overview of our CustomStudio dataset. Samples in our CustomStudio dataset cover a wide range of object categories.